OpenStudy (anonymous):

Ten persons sit around a table for a dinner. The chance that two specified persons are not sitting side by side, is

OpenStudy (anonymous):

9! 2 perons so 7!

OpenStudy (anonymous):

7!/9!

OpenStudy (mokeira):

9-8!/9! 7!/9! you are right

OpenStudy (anonymous):

how 9-8!/9! is 7!/9!

OpenStudy (anonymous):

how @Mokeira

OpenStudy (mokeira):

my bad! lol

OpenStudy (mokeira):

sorry about that..there are 10 people so (8-1)!/9! that is what i meant so that you get 7!/9!

OpenStudy (anonymous):

according to the question, no. of ways in which no 2 persons sit together = (total no. of ways in which 10 persons can be arranged around the circular table) minus (no. of ways in which 2persons sit together) so, no. of ways in which 10 persons can be arranged= (10-1)!=9! no. of ways if 2 persons sit together= (9-1)!*2= 8!*2 hence, no. of ways in which no 2 persons sit together =9!-{8!*2}= 9*8!-{8!*2} =8!{9-2}=8!*7= 282240